there are 10x+5 chairs in the auditorium.if there are 2x+1 chairs in each row,how many rows are there?

Answer by Cromlix(4381)

Hi there,
Make the number of rows = x
Rows x Chairs = total number of chairs.
x(2x + 1) = 10x + 5
2x^2 + x = 10x + 5
Put in ax^2 + bx + c form
2x^2 - 9x - 5 = 0
(2x + 1)(x - 5) = 0
2x + 1 = 0 x = -1/2 (Disregard as -ve)
x - 5 = 0
x = 5
There are 5 rows in the auditorium.
Hope this helps :-)